Clyde, a stubborn 380 lbs mule, refuses to walk into the horse trailer, so Farmer Brown...

  1. Clyde, a stubborn 380 lbs mule, refuses to walk into the horse trailer, so Farmer Brown must drag him up a 15 ft ramp, which stands 1.5 ft above the ground.

A) What is the ideal mechanical advantage of the ramp?

B) If Farmer Brown needs to exert a force of 60 lbs on the mule, what is the actual mechanical advantage of the ramp?

C) What is the efficiency of the ramp?
Answer #1

a)

The Ideal Mechanical advantage of the ramp is

IMA=\frac{15}{1.5}=10

b)

The actual mechanical advantage of the ramp is

AMA=\frac{380}{60}=6.33\simeq 6.3

c)

The efficiency of the ramp is

\eta =\frac{6.33}{10}\times 100=63.3 %